OpenStudy (precal):

Find the derivative of y=log(4x)

OpenStudy (anonymous):

hint : log(4x) = ln(4x)/ln(10)

OpenStudy (precal):

\[\frac{ d }{ dx }\left( \log _{a} (x)\right)=\frac{ 1 }{ xln(a) }\]

OpenStudy (anonymous):

correct

OpenStudy (precal):

yes, you used change of base

OpenStudy (anonymous):

it is better to change base than remember this formula. if you change base you only need to know the derivative of ln

OpenStudy (precal):

yes that would be easier let me try it on paper.

OpenStudy (precal):

then you used quotient rule correct?

OpenStudy (anonymous):

no, ln(10) is just a number :)

OpenStudy (precal):

you lost me

OpenStudy (anonymous):

taking the derivative of ln(4x)/ln(10) is not requiring quotient rule since ln(10) is a number and hence you dont have to differentiate it

OpenStudy (precal):

ok so my answer is u' over u correct?

OpenStudy (precal):

ok got it, u=4x u'=4

OpenStudy (precal):

using the formula above it is \[\frac{ 4 }{ 4x(\ln10) }\]

OpenStudy (precal):

4/4 is one so my solution is 1/(x(ln10)) is your way faster

OpenStudy (precal):

ok got it, I used the formula and the chain rule. My solutions match the key....

OpenStudy (anonymous):

you can always use this formula but then you have to remember it. i prefer to change base